A stone is vertically thrown upward with a speed of 13 m/s from the edge of a cliff 60m high.how much later does it reach the bottom of the cliff
1
Expert's answer
2021-08-02T08:41:02-0400

The kinematic equation of motion gives the following (y-axis directed upward, origin is at ground):


"y = y_0 + v_0t - \dfrac{gt^2}{2}"

where "y_0 = 60m" is the inital coordinate (cliff height), "v_0 = 13m\/s" is the inital speed, "g = 9.81m\/s^2" is the gravitational acceleration, and "y" is the coordinate at time "t".

At the bottom of the cliff "y = 0", this, substituting it, obtain:


"0 = 60 + 13t - 4.905t^2\\
4.905t^2 - 13t - 60 = 0"

Solving the quadratic equation and taking the positive solution, obtain:


"t \approx 7.7s"

Answer. 7.7 s.